OpenStudy (chrisneedshelp):

Find the exact value of the following. Do NOT use labels in your answer and do NOT round your answer. Find -xy, for x = -75.9 and y = 2.6. -xy =

OpenStudy (anonymous):

9867/50 as a fraction

OpenStudy (chrisneedshelp):

hm?

OpenStudy (anonymous):

to get an exact value, convert the answer into a fraction

OpenStudy (anonymous):

x = -759/10 y = 26/10 -x = 759/10 -xy = 759/10 x 26/10 -xy = 19734/100 - xy = 9867/50

OpenStudy (anonymous):

-xy=-(-75.9*2.6)= 197.34

OpenStudy (anonymous):

^^^ that is not exact form

OpenStudy (anonymous):

it doesn't look like he understood what you did. so maybe he is on a lower class then you thought. but yeah your method is the correct one if he wants exact form

OpenStudy (anonymous):

yeah so to get an exact value, simply multiply 197.34 by 100 and you get 19734/100. simplify and you will get 9867/50